# X-Adapter: Adding Universal Compatibility of Plugins for Upgraded Diffusion Model

### v2

@@ Abstract
Figure 1: Given pretrained plug-and-play modules (e.g., ControlNet, LoRA) of the base diffusion model (e.g., Stable Diffusion 1.5), the proposed X-Adapter can universally upgrade these plugins, enabling them directly work with the upgraded Model (e.g., SDXL) without further retraining. Text prompts: “1girl, solo, smile, looking at viewer, holding flowers” “Apply face paint” “1girl, upper body, flowers” “A cute cat holding a gun” “Best quality, extremely detailed” “A fox made of water” from left to right, top to bottom.
@@ 5 Conclusion
In this paper, we target a new task of upgrading all the downstream plugins trained on old diffusion model to the upgraded ones. To this end, we propose X-Adapter, which comprises a copied network structure and weights of the base model, and a series of mapping layers between two decoders for feature mapping. During training, we freeze the upgraded model and set its text prompt to empty text to maximize the function of X-Adapter. In testing, we propose a two-stage inference strategy to further enhance performance. We conduct comprehensive experiments to demonstrate the advantages of the proposed methods in terms of compatibility and visual quality.
Supplementary Material
Appendix A Ablation on module selection
We provide quantitative results of ablation on module to insert mapping layers. Tab. 1 indicates that the decoder-only strategy shows the strongest guidance capability since it does not harm the encoder’s feature space and only performs guidance during generation.
Appendix B Limitation
For plugins to generate personalized concepts, e.g., IP-Adapter [39], our method might not maintain the identity well like shown in Fig. 1. This is because the custom plugins work on the text-encoder other than the feature space concepts that are not directly injected into the upgraded model other than fused as guidance.
Appendix C Qualitative result
We provide more qualitative result on Controlnet [43], T2I-Adapter [24] and LoRA [17] as shown in Fig. 2
Encoder & Decoder 38.37 0.26 0.24 ± 0.15
Encoder only 37.32 0.26 0.23 ± 0.14
Decoder only 30.95 0.26 0.27 ± 0.13
Plugin: LoRA FID $\downarrow$ CLIP-score $\uparrow$ Style-Sim $\uparrow$
Encoder & Decoder 36.71 0.26 0.79
Encoder only 35.54 0.26 0.80
Decoder only 29.88 0.26 0.83
Table 1: Ablation on module to insert mapping layers. For encoder only and decoder only, we use the same mapping layers to map the frozen base model to the upgraded model’s encoder and decoder separately. For both the encoder and decoder, we use two identical mapping layers and insert them into the encoder and decoder.
Figure 1: Limitations. In IP-Adapter [39], although our method can produce relatively identity-consistent results, the details, e.g., clothes, are still different from the original model.
Figure 2: Qualitative Results on Different Plugins.

### v3

@@ Abstract
Figure 1: Given pretrained plug-and-play modules (e.g., ControlNet, LoRA) of the base diffusion model (e.g., Stable Diffusion 1.5), the proposed X-Adapter can universally upgrade these plugins, enabling them directly work with the upgraded Model (e.g., SDXL) without further retraining. Text prompts: “1girl, solo, smile, looking at viewer, holding flowers” “Apply face paint” “1girl, upper body, flowers” “A colorful lotus, ink” “Best quality, extremely detailed” “A fox made of water” from left to right, top to bottom.
@@ 5 Conclusion
In this paper, we target a new task of upgrading all the downstream plugins trained on old diffusion model to the upgraded ones. To this end, we propose X-Adapter, which comprises a copied base model and mapping layers between two decoders for feature mapping. During training, we freeze the upgraded model and set its text prompt to empty text to maximize the function of X-Adapter. In testing, we propose a two-stage inference strategy to further enhance performance. We conduct comprehensive experiments to demonstrate the advantages of the proposed methods.
6 Acknowledge
This research is supported by National Research Foundation, Singapore and A*STAR, under its RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) grant call (Grant No. I2001E0059) – SIA-NUS Digital Aviation Corp Lab. Mike Zheng Shou is supported by the Ministry of Education, Singapore, under the Academic Research Fund Tier 1 (FY2022) Award 22-5406-A0001.